# BaB-ND: Long-Horizon Motion Planning with Branch-and-Bound and Neural Dynamics

### v2

@@ 1 Introduction
To address these challenges, the community has developed a range of approaches. Sampling-based methods such as the Cross-Entropy Method (CEM) (Rubinstein & Kroese, 2013) and Model Predictive Path Integral (MPPI) (Williams et al., 2017) have gained popularity in manipulation tasks (Lowrey et al., 2018; Manuelli et al., 2020; Nagabandi et al., 2020; Wang et al., 2023) due to their flexibility, compatibility with neural dynamics models, and strong GPU support. However, their performance in more complex, higher-dimensional planning problems is limited and requires further theoretical analysis (Yi et al., 2024). Alternatively, more principled optimization approaches, such as Mixed-Integer Programming (MIP), have been applied to planning problems using sparsified neural dynamics models with ReLU activations (Liu et al., 2023). Despite achieving global optimality and better closed-loop control performance, MIP is inefficient and struggles to scale to large neural networks, limiting its ability to handle larger-scale planning problems.
@@ 2 Related Works
The highly non-linear and non-convex nature of neural dynamics models hinders the effective optimization of model-based planning problems. Previous works (Yen-Chen et al., 2020; Ebert et al., 2017; Nagabandi et al., 2020; Finn & Levine, 2017; Manuelli et al., 2020; Sacks et al., 2023; Han et al., 2024) utilize sampling-based algorithms like CEM (Rubinstein & Kroese, 2013) and MPPI (Williams et al., 2017) for online planning. Despite their flexibility and ability to leverage GPU support, these methods struggle with large input dimensions due to the exponential growth in the number of required samples. Previous work (Yin et al., 2022) improved MPPI by introducing dynamics model linearization and covariance control techniques, but their effectiveness when applied to neural dynamics models remains unclear. Other approaches (Li et al., 2018; 2019) have used gradient descent to optimize action sequences but encounter challenges with the local optima and non-smooth objective landscapes. Recently, methods inspired by neural network verification have been developed to achieve safe control and robust planning over systems involving neural networks (Wei & Liu, 2022; Liu et al., 2023; Hu et al., 2024a; Wu et al., 2024; Hu et al., 2024b), but their scalability to more complex real-world manipulation tasks is still uncertain. Moreover, researchers are also exploring the promising direction of performing planning over graphs of convex sets (GCSs) for contact-rich manipulation tasks Marcucci (2024); Graesdal et al. (2024). However, these approaches do not incorporate neural networks.
@@ 3 Branch-and-Bound for Planning With Neural Dynamics Models
We formulate the planning problem as an optimization problem in  Eq. 1, where $c$ is the cost function, $t_{0}$ is the current time step, and $H$ is the planning horizon. $\hat{x}_{t}$ is the (predicted) state at time step $t$, and the current state $\hat{x}_{t_{0}}=x_{t_{0}}$ is known. $u_{t}\in\{u\mid\underline{u}\leq u\leq\overline{u}\}\subset\mathbb{R}^{k}$ is the robot’s action at each step. $f_{dyn}$ is the pre-trained neural dynamics model (Please refer to Section D.3 for details about learning the neural dynamics model.), which takes state and action at time $t$ and predicts the next state $\hat{x}_{t+1}$. The goal of the planning problem is to find a sequence of optimal actions $u_{t}$ that minimize the sum of step costs:
$\min_{\{u_{t}\in{\mathcal{U}}\}}\sum_{t=t_{0}}^{t_{0}+H}c(\hat{x}_{t},u_{t})\quad\textrm{s.t.}\quad\hat{x}_{t+1}=f_{dyn}(\hat{x}_{t},u_{t})\end{aligned}\qquad\Longrightarrow\qquad\min_{{\bm{u}}\in{\mathcal{C}}}f({\bm{u}})$ (1)
This problem can be challenging due to its long planning horizon $H$, complex cost function $c$, and the non-linear neural dynamics model $f_{dyn}$ with recursive dependencies at every step. Existing sampling-based and gradient-based methods may converge to sub-optima without systematic searching, while MIP-based methods fail to scale up with the size of $f_{dyn}$ and the planning horizon $H$.
To simplify notation, we can substitute all constraints on $\hat{x}_{t+1}$ into the objective recursively, and further simplify the problem as a constrained optimization problem $\min_{{\bm{u}}\in{\mathcal{C}}}f({\bm{u}})$ (Eq. 1). Here $f$ is our final objective, a scalar function that absorbs the neural network $f_{dyn}$ and the cost function summed in all $H$ steps. ${\bm{u}}=\{u_{t_{0}:t_{0}+H}\}\in{\mathcal{C}}$ is the action sequence and ${\mathcal{C}}\subset\mathbb{R}^{d}$ is the entire input space with dimension $d=kH$. We also flatten ${\bm{u}}$ as a vector containing actions for all time steps, and use ${\bm{u}}_{j}$ to denote a specific dimension. Our goal is to then find the optimal objective value $f^{*}$ and its corresponding optimal action sequence ${\bm{u}}^{*}$.
@@
Although this generic BaB framework has been used in neural network verifiers (Bunel et al., 2018; Wang et al., 2021), to prove a sound lower bound of $f({\bm{u}})$ within $\mathcal{C}$, $\underline{f}^{*}_{\mathcal{C}}\leq f^{*}$, whereas our BaB-ND seeks a concrete solution $\tilde{{\bm{u}}}$ (a near-optimal action sequence) to an objective-minimization problem $\min_{{\bm{u}}\in\mathcal{C}}{f({\bm{u}})}$. These fundamental distinctions in goals lead to different design choices.
@@ 3.1 Branching Heuristics for BaB-ND Planning
The efficiency of BaB heavily depends on the quality of branches. Hence, selecting promising subdomains and determining how to split subdomains are two essential questions in BaB, referring to $\mathtt{batch\_pick\_out}({\mathbb{P}},n)$ and $\mathtt{batch\_split}\left(\{{\mathcal{C}}_{i}\}\right)$ in Algorithm 1. Here we introduce our specialized branching heuristics to select and split subdomains for seeking high-quality solutions.
@@
The function $\mathtt{batch\_pick\_out}({\mathbb{P}},n)$ picks $n$ most promising subdomains for branching, based on their associated $\underline{f}^{*}_{{\mathcal{C}}_{i}}$ or $\overline{f}^{*}_{{\mathcal{C}}_{i}}$. The pickout process must balance exploitation (focusing on areas around good solutions) and exploration (investigating regions that have not been thoroughly explored). First, we sort subdomains ${\mathcal{C}}_{i}$ by $\overline{f}^{*}_{{\mathcal{C}}_{i}}$ in ascending order and select the first $n_{1}$ subdomains to form $\{{\mathcal{C}}^{1}_{\text{pick}}\}$. Subdomains with smaller $\overline{f}^{*}_{{\mathcal{C}}_{i}}$ are prioritized, as good solutions have been found there. Then, we form another promising set $\{{\mathcal{C}}^{2}_{\text{pick}}\}$ by sampling $n-n_{1}$ subdomains from the remaining $N$ ones, using $\mathrm{softmax}$, with the probability $p_{i}$ defined in Eq. 2, where $T$ is the temperature and $\underline{f}^{*}_{{\mathcal{C}}_{i},\text{scaled}}$ represents $\underline{f}^{*}_{{\mathcal{C}}_{i}}$ after min-max normalization for numerical stability. A smaller $\underline{f}^{*}_{{\mathcal{C}}_{i}}$ may indicate potentially better solutions in ${\mathcal{C}}_{i}$, which should be prioritized.
$p_{i}=\frac{\exp(-\underline{f}^{*}_{{\mathcal{C}}_{i},\text{scaled}}/T)}{\sum_{j=1}^{N}\exp(-\underline{f}^{*}_{{\mathcal{C}}_{j},\text{scaled}}/T)}$ (2)
@@ 4 Experimental Results
In this section, we assess the performance of our BaB-ND across a variety of complex robotic manipulation tasks. Our primary objective is to address three key questions through experiments: 1) How effectively does our BaB-ND perform long-horizon planning? 2) Is our BaB-ND applicable to different manipulation scenarios with multi-object interactions and deformable objects? 3) What is the scalability of our BaB-ND compared to existing methods?
@@
In Figure 4, we visualize the best objective values found by different methods across different input dimensions up to $d=100$. BaB-ND consistently outperforms all baselines which converge to sub-optimal values. For $d=100$, BaB-ND can achieve optimality on 98 to 100 dimensions. This synthetic experiment demonstrates the potential of BaB-ND on planning problems involving neural dynamics.
@@
  • Pushing with Obstacles. In Figure 5.a, this task involves using a pusher to manipulate a “T”-shaped object to reach a target pose while avoiding collisions with obstacles. An MLP neural dynamics model is trained with interactions between the pusher and object without obstacles. Obstacles are modeled in the cost function, making non-smooth landscapes and non-convex feasible regions.
@@
Figure 5: Qualitative results on real-world manipulation tasks. We evaluate our BaB-ND across four complex robotic manipulation tasks, involving non-convex feasible regions, requiring long-horizon planning, and interactions between multiple objects and the deformable rope. For each task, we visualize the initial and target configurations and one successful trajectory. Please refer to our project page for demonstrations.
@@
Figure 6: Quantitative analysis of planning performance and execution performance in real world. BaB-ND consistently outperforms baselines on open-loop performance leading to better closed-loop performance. (a) The open-loop performance of all tasks in simulation. We report the best objective of Eq. 1 found by all methods across different test cases. (b) The closed-loop performance in real world. GD is excluded from testing due to its poor open-loop performance. We report the success rate for Rope Routing, since a greedy trajectory that horizontally routes the rope may achieve a low final cost but fails to place it into the slot, and we report final step costs for the other tasks.
@@
In the Rope Routing task, MPPI, CEM and ours achieve similar open-loop performance while GD may struggle at sub-optimal trajectories, routing the rope horizontally and becoming stuck outside the slot. In the Object Sorting task, CEM can outperform MPPI in simulation and the real world since MPPI is more suitable for planning continuous action sequences while actions are discrete in the task. Ours outperforms CEM with similar median and smaller variance.
@@
We evaluate the scalability of our BaB-ND on Pushing with Obstacles task with varying model sizes and planning horizons on multiple test cases, in comparison to MIP (Liu et al., 2023). We train the neural dynamics models with different sizes and the same architecture, and use the number of parameters in the single neural dynamics model $f_{dyn}$ to indicate the model size. Moreover, to accommodate MIP, we remove all obstacle-related components in cost and define the objective as the step cost after planning horizon $H$, $c(s_{t_{0}+T},x_{t_{0}+T},s_{\text{goal}})$ instead of the accumulated cost.
In Figure 7 (a), we visualize the average runtime of MIP and ours on test cases with different model sizes and planning horizons. The results show that MIP only handles small problems. Among all 36 settings, it provides optimal solutions for 6, sub-optimal solutions for 3, and fails to find any solution for the remaining within 300 seconds. In contrast, our BaB-ND scales up well to large problems with planning horizon $H=20$ and a model containing over 500K parameters.
In Figure 7 (b), we evaluate the runtime of each primary component of our BaB-ND across various model sizes, ranging from approximately 9K to over 500K, in the context of an original objective for the Pushing with Obstacles tasks (containing items to model obstacles and accumulated cost among all steps) over a planning horizon of $H=20$. The breakdown bar chart illustrates that the runtimes for the branching and bounding components grow relatively slowly across model sizes, which increase more than 50-fold. Our improved bounding procedure, as discussed in Section 3.2, scales well with growing model size. In addition, the searching runtime scales proportionally to neural network size since the majority of searching time is spent on sampling the model with a large batch size on GPUs.
@@ 6 Acknowledgment
This work is supported by Toyota Research Institute (TRI). We thank Jose Barreiros, Hongkai Dai, Aykut Onol and Mengchao Zhang for their constructive suggestions on the paper manuscript. We also thank Mingtong Zhang, Haozhe Chen, Baoyu Li and Binghao Huang for their help with real-world experiments and simulation environment development. This article reflects solely the opinions and conclusions of its authors, not those of TRI or any other Toyota entity.
@@ C.1 Scalability analysis
We conducted an experiment to compare the scalability of our BaB-ND with sampling-based methods on complex planning problems. We used the same model sizes and planning horizons as in Figure 7 (a), optimizing the complex objective function applied in the Pushing with Obstacles task. Parameters for all methods were adjusted to ensure similar runtimes for the largest problems.
@@ C.3 Ablation Study and Hyper-Parameter Analysis
We conduct an additional ablation study on the Pushing with Obstacles and Object Sorting tasks to evaluate how different design choices impact planning performance in Table 4.
@@
We adjust three hyper-parameters in BaB-ND for the tasks Pushing with Obstacles and Object Sorting to evaluate its hyper-parameter sensitivity:

### v3

@@ 1 Introduction
To address these challenges, the community has developed a range of approaches. Sampling-based methods such as the Cross-Entropy Method (CEM) (Rubinstein & Kroese, 2013) and Model Predictive Path Integral (MPPI) (Williams et al., 2017) have gained popularity in manipulation tasks (Lowrey et al., 2018; Manuelli et al., 2020; Nagabandi et al., 2020; Wang et al., 2023) due to their flexibility, compatibility with neural dynamics models, and strong GPU support. However, their performance in more complex, higher-dimensional planning problems is limited and requires further theoretical analysis (Yi et al., 2024). Alternatively, more principled optimization approaches, such as Mixed-Integer Programming (MIP), have been applied to planning problems using sparsified neural dynamics models with ReLU activations (Liu et al., 2023b). Despite achieving global optimality and better closed-loop control performance, MIP is inefficient and struggles to scale to large neural networks, limiting its ability to handle larger-scale planning problems.
@@ 2 Related Works
The highly non-linear and non-convex nature of neural dynamics models hinders the effective optimization of model-based planning problems. Previous works (Yen-Chen et al., 2020; Ebert et al., 2017; Nagabandi et al., 2020; Finn & Levine, 2017; Manuelli et al., 2020; Sacks et al., 2023; Han et al., 2024) utilize sampling-based algorithms like CEM (Rubinstein & Kroese, 2013) and MPPI (Williams et al., 2017) for online planning. Despite their flexibility and ability to leverage GPU support, these methods struggle with large input dimensions due to the exponential growth in the number of required samples. Previous work (Yin et al., 2022) improved MPPI by introducing dynamics model linearization and covariance control techniques, but their effectiveness when applied to neural dynamics models remains unclear. Other approaches (Li et al., 2018; 2019) have used gradient descent to optimize action sequences but encounter challenges with the local optima and non-smooth objective landscapes. Recently, methods inspired by neural network verification have been developed to achieve safe control and robust planning over systems involving neural networks (Wei & Liu, 2022; Liu et al., 2023b; Hu et al., 2024a; Wu et al., 2024; Hu et al., 2024b), but their scalability to more complex real-world manipulation tasks is still uncertain. Moreover, researchers are also exploring the promising direction of performing planning over graphs of convex sets (GCSs) for contact-rich manipulation tasks Marcucci (2024); Graesdal et al. (2024). However, these approaches do not incorporate neural networks.
@@ 3 Branch-and-Bound for Planning With Neural Dynamics Models
We formulate the planning problem as an optimization problem in  Eq. 1, where $c$ is the cost function, $t_{0}$ is the current time step, and $H$ is the planning horizon. $\hat{x}_{t}$ is the (predicted) state at time step $t$, and the current state $\hat{x}_{t_{0}}=x_{t_{0}}$ is known. $u_{t}\in\{u\mid\underline{u}\leq u\leq\overline{u}\}\subset\mathbb{R}^{k}$ is the robot’s action at each step. $f_{\text{dyn}}$ is the pre-trained neural dynamics model (Please refer to Section D.3 for details about learning the neural dynamics model.), which takes state and action at time $t$ and predicts the next state $\hat{x}_{t+1}$. The goal of the planning problem is to find a sequence of optimal actions $u_{t}$ that minimize the sum of step costs:
$\min_{\{u_{t}\in{\mathcal{U}}\}}\sum_{t=t_{0}}^{t_{0}+H}c(\hat{x}_{t},u_{t})\quad\textrm{s.t.}\quad\hat{x}_{t+1}=f_{\text{dyn}}(\hat{x}_{t},u_{t})\end{aligned}\qquad\Longrightarrow\qquad\min_{{\bm{u}}\in{\mathcal{C}}}f({\bm{u}}).$ (1)
This problem can be challenging due to its long planning horizon $H$, complex cost function $c$, and the non-linear neural dynamics model $f_{\text{dyn}}$ with recursive dependencies at every step. Existing sampling-based and gradient-based methods may converge to sub-optima without systematic searching, while MIP-based methods fail to scale up with the size of $f_{\text{dyn}}$ and the planning horizon $H$.
To simplify notation, we can substitute all constraints on $\hat{x}_{t+1}$ into the objective recursively, and further simplify the problem as a constrained optimization problem $\min_{{\bm{u}}\in{\mathcal{C}}}f({\bm{u}})$ (Eq. 1). Here $f$ is our final objective, a scalar function that absorbs the neural network $f_{\text{dyn}}$ and the cost function summed in all $H$ steps. ${\bm{u}}=\{u_{t_{0}:t_{0}+H}\}\in{\mathcal{C}}$ is the action sequence and ${\mathcal{C}}\subset\mathbb{R}^{d}$ is the entire input space with dimension $d=kH$. We also flatten ${\bm{u}}$ as a vector containing actions for all time steps, and use ${\bm{u}}_{j}$ to denote a specific dimension. Our goal is to then find the optimal objective value $f^{*}$ and its corresponding optimal action sequence ${\bm{u}}^{*}$.
@@
Although the generic BaB framework has been applied in neural network verification (Bunel et al., 2018; Wang et al., 2021), its goal is to prove a sound lower bound of $f({\bm{u}})$ within $\mathcal{C}$, ensuring that $\underline{f}^{*}_{\mathcal{C}}\leq f^{*}$. In contrast, our BaB-ND aims to find a concrete solution $\tilde{{\bm{u}}}$—a near-optimal action sequence—for the objective-minimization problem $\min_{{\bm{u}}\in\mathcal{C}}f({\bm{u}})$. These fundamental differences in objectives lead to distinct design choices.
@@ 3.1 Branching Heuristics for BaB-ND Planning
The efficiency of BaB relies heavily on the quality of branching. Thus, selecting promising subdomains and determining effective ways to split them are two critical aspects of BaB, corresponding to $\mathtt{batch\_pick\_out}({\mathbb{P}},n)$ and $\mathtt{batch\_split}\left(\{{\mathcal{C}}_{i}\}\right)$ in Algorithm 1. Here, we introduce our specialized branching heuristics designed to select and split subdomains to obtain high-quality solutions.
@@
The function $\mathtt{batch\_pick\_out}({\mathbb{P}},n)$ picks $n$ most promising subdomains for branching, based on their associated $\underline{f}^{*}_{{\mathcal{C}}_{i}}$ or $\overline{f}^{*}_{{\mathcal{C}}_{i}}$. The pickout process must balance exploitation (focusing on areas around good solutions) and exploration (investigating regions that have not been thoroughly explored). First, we sort subdomains ${\mathcal{C}}_{i}$ by $\overline{f}^{*}_{{\mathcal{C}}_{i}}$ in ascending order and select the first $n_{1}$ subdomains to form $\{{\mathcal{C}}^{1}_{\text{pick}}\}$. Subdomains with smaller $\overline{f}^{*}_{{\mathcal{C}}_{i}}$ are prioritized, as good solutions have been found there. Then, we form another promising set $\{{\mathcal{C}}^{2}_{\text{pick}}\}$ by sampling $n-n_{1}$ subdomains from the remaining $N$ ones, using $\mathrm{softmax}$, with the probability $p_{i}$ defined in Eq. 2, where $T$ is the temperature and $\underline{f}^{*}_{{\mathcal{C}}_{i},\text{scaled}}$ represents $\underline{f}^{*}_{{\mathcal{C}}_{i}}$ after min-max normalization for numerical stability. A smaller $\underline{f}^{*}_{{\mathcal{C}}_{i}}$ may indicate potentially better solutions in ${\mathcal{C}}_{i}$, which should be prioritized:
$p_{i}=\frac{\exp(-\underline{f}^{*}_{{\mathcal{C}}_{i},\text{scaled}}/T)}{\sum_{j=1}^{N}\exp(-\underline{f}^{*}_{{\mathcal{C}}_{j},\text{scaled}}/T)}.$ (2)
@@ 4 Experimental Results
In this section, we evaluate the performance of our BaB-ND on a range of complex robotic manipulation tasks. Our primary objective is to investigate three key questions through experiments: (1) How effectively does BaB-ND handle long-horizon planning? (2) Is BaB-ND applicable to diverse manipulation scenarios involving multi-object interactions and deformable objects? (3) How does the scalability of BaB-ND compare to existing methods?
@@
In Figure 4, we visualize the best objective values found by different methods across different input dimensions up to $d=100$. BaB-ND consistently outperforms all baselines which converge to sub-optimal values. For $d=100$, BaB-ND can achieve optimality on 98 to 100 dimensions. This synthetic experiment demonstrates the potential of BaB-ND in planning problems involving neural dynamics.
@@
  • Pushing with Obstacles. In Figure 5.a, this task involves using a pusher to manipulate a “T”-shaped object to reach a target pose while avoiding collisions with obstacles. An MLP neural dynamics model is trained with interactions between the pusher and object without obstacles. Obstacles are modeled in the cost function, resulting in non-smooth landscapes and non-convex feasible regions.
@@
Figure 5: Qualitative results on real-world manipulation tasks. We evaluate our BaB-ND across four complex robotic manipulation tasks, involving non-convex feasible regions, requiring long-horizon planning, and interactions between multiple objects and the deformable rope. For each task, we visualize the initial and target configurations and one successful trajectory. Please refer to our project page for video demonstrations.
@@
Figure 6: Quantitative analysis of planning performance and execution performance in real world. BaB-ND consistently outperforms baselines on open-loop performance leading to better closed-loop performance. (a) The open-loop performance of all tasks in simulation. We report the best objective of Eq. 1 found by all methods across different test cases. (b) The closed-loop performance in real world. GD is excluded from testing due to its poor open-loop performance. We report the success rate for Rope Routing, since a greedy trajectory that horizontally routes the rope may achieve a low final cost but fails to place it into the slot, and we report final step costs for all other tasks.
@@
In the Rope Routing task, MPPI, CEM, and our method achieve comparable open-loop performance, while GD may struggle with suboptimal trajectories, often routing the rope horizontally and getting stuck outside the slot. In the Object Sorting task, CEM outperforms MPPI, as MPPI is better suited for planning continuous action sequences, whereas this task involves discrete actions. Our method outperforms CEM, achieving a similar median performance with noticeably lower variance.
@@
We evaluate the scalability of our BaB-ND on Pushing with Obstacles task with varying model sizes and planning horizons on multiple test cases, in comparison to MIP (Liu et al., 2023b). We train the neural dynamics models with different sizes and the same architecture, and use the number of parameters in the single neural dynamics model $f_{\text{dyn}}$ to indicate the model size. Moreover, to accommodate MIP, we remove all obstacle-related components in cost and define the objective as the step cost after planning horizon $H$, $c(\hat{x}_{t_{0}+H},u_{t_{0}+H})$ instead of the accumulated cost.
In Figure 7.a, we visualize the average runtime of MIP and ours on test cases with different model sizes and planning horizons. The results show that MIP only handles small problems. Among all 36 settings, it provides optimal solutions for 6, sub-optimal solutions for 3, and fails to find any solution for the remaining within 300 seconds. In contrast, our BaB-ND scales up well to large problems with planning horizon $H=20$ and a model containing over 500K parameters.
In Figure 7.b, we evaluate the runtime of each primary component of our BaB-ND across various model sizes, ranging from approximately 9K to over 500K, using the original objective for the Pushing with Obstacles tasks (containing items to model obstacles and accumulated cost among all steps) over a planning horizon of $H=20$. The breakdown bar chart illustrates that the runtimes for the branching and bounding components grow relatively slowly across model sizes, which increase more than 50-fold. Our improved bounding procedure, as discussed in Section 3.2, scales well with growing model size. In addition, the searching runtime scales proportionally to neural network size since the majority of searching time is spent on sampling the model with a large batch size on GPUs.
@@ 6 Acknowledgment
This work is supported by the Toyota Research Institute (TRI). We thank Hongkai Dai, Aykut Onol, and Mengchao Zhang for their constructive suggestions on the paper manuscript. We also thank Mingtong Zhang, Haozhe Chen, Baoyu Li, and Binghao Huang for their help with real-world experiments and simulation environment development. This article reflects solely the opinions and conclusions of its authors, not those of TRI or any other Toyota entity.
@@ C.1 Scalability analysis
We conducted an experiment to compare the scalability of our BaB-ND with sampling-based methods on complex planning problems. We used the same model sizes and planning horizons as in Figure 7 (a), optimizing the complex objective function applied in the Pushing w/ Obstacles task. Parameters for all methods were adjusted to ensure similar runtimes for the largest problems.
@@ C.3 Ablation Study and Hyper-Parameter Analysis
We conduct an additional ablation study on the Pushing w/ Obstacles and Object Sorting tasks to evaluate how different design choices impact planning performance in Table 4.
@@
We adjust three hyper-parameters in BaB-ND for the tasks Pushing w/ Obstacles and Object Sorting to evaluate its hyper-parameter sensitivity: